equation:
sqrt(x+7) - sqrt(x+4)= 1
solve for x.

Answer by stanbon!(97)

Isolate one of the radicals on the left side of the equal sign as follows:
sqrt(x+7) = sqrt(x+4) + 1
Square both sides to get:
x+7 = x+4 + 2sqrt(x+4) +1
Combine like terms and isolate the remaining sqrt term as follows:
2sqrt(x+4)=2
Simplify to get:
sqrt(x+4) = 1
Square both sides to get:
x+4 = 1
x = -3
Check the answer in the original equation because
all this squaring can produce extraneous answers.
Cheers,
Stan H.